Calculate the resultant of two rectangular simple harmonic vibration of whose amplitude as well as periods are in the ratio 1:2 and the phase difference is 90 degress? please give solution and send my email id mynameisswamy26@gmail.com

Grade:12

2 Answers

Chetan Mandayam Nayakar
13 years ago

x=acos(2wt), y= 2acos(wt+(π/2)) = -2asin(wt)

From inspection, it is obvious that resultant T=2π/w, let A=resultant amplitude. A2= max(a2(4sin2(wt)+(1-2sin2(wt))2)) =5a2

thus, A =a√5

amit yadav
7 years ago
calculat the wavelength of longitudinal waves of frenquecy 600hz in asolid whose density is 6.5×103kg/m3 and young ‘s modulus 8.0×1010newton/Meter2